OpenStudy (anonymous):

a parallelogram with sides of 6 and 10 has an area of 30 square root 2.find the measure of the small angle of the parallelogram.

OpenStudy (anonymous):

area of parallelogram=base * height so,10*h=30 h=3 so from the figure.... sin(o)=3/6=1/2... o=30* this is the required ans...

OpenStudy (anonymous):

how did yu find 3/6=1/2? do you know how to find square roots like this: 4

OpenStudy (anonymous):



OpenStudy (anonymous):

hey,its quite simple......